x^2 + 3x - 18 = 0
Use the following quadratic equation to complete parts a and b below. You must show all work.
a. Solve by factoring.
b. Solve using the quadratic formula

Answer by awang1996(50)

x^2 + 3x - 18 = 0
(x+6)(x-3)=0
x=-6 x=3
________________________________________________________________________________

quadratic formula:
-b plus or minus the square root of "b^2-4ac"
----------------------------------------------
2a
x^2 + 3x - 18 = 0

-3 plus or minus the square root of "(-3)^2-4(1)(-18)
-----------------------------------------------------
2(1)

-3 plus or minus the square root of "9+72"
-------------------------------------------
2

-3 plus or minus the square root of "81"
-----------------------------------------
2

-3 plus or minus "9"
--------------------
2

the two equations you have:
-3+9 -3-9
---- -----
2 2

6 -12
- ----
2 2

3 -6


x={3,-6}